OpenStudy (anonymous):

How to find all real or imaginary zeros for each polynomial function: F (x) = 2x^3-4x^2+3x F (x)= x^3+2x^2-6x I know how to do these type of problems but, both do not have a solution found in the given P/Q set.... so, I do not know what to do... help! (This was not covered in lecture & exam is tomorrow !)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle f(x)=2x^3-4x^2+3x }\)

OpenStudy (solomonzelman):

To find the zero, set: \(\large\color{black}{ \displaystyle 0=2x^3-4x^2+3x }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 0=x(2x^2-4x+3) }\)

OpenStudy (solomonzelman):

So you have two cases: [Case 1] \(\large\color{black}{ \displaystyle x=0 }\) [Case 2] \(\large\color{black}{ \displaystyle x=2x^2-4x+3 }\)

OpenStudy (anonymous):

Ok am following so far....

OpenStudy (solomonzelman):

Ok, the first case is already a solution, and case 2 can be solved using the quadratic formula.

OpenStudy (anonymous):

So we, factor further right ? (The second case)

OpenStudy (triciaal):

corrrect

OpenStudy (anonymous):

Ok let me try......

OpenStudy (solomonzelman):

You can't really factor the second case, but you can use the quadratic formula.

OpenStudy (solomonzelman):

\(\bbox[8pt, yellow ,border:8px solid black]{\Huge \Huge{x=~} \Huge{ \frac{-\color{magenta}{b} \pm\sqrt{ \color{magenta}{b} ^2-4 \color{blue}{a} \color{red}{c}}}{2 \color{blue}{a}} } ~ }\) when the equation is \(\LARGE\color{black}{ \color{blue}{a} x^2+ \color{magenta}{b}x+ \color{red}{c}=0 }\).

OpenStudy (anonymous):

Oh, I remember now.... ok let me try that! Once I sove I have my answer ?

OpenStudy (anonymous):

Ahh.... I think am doing it wrong... well I have

OpenStudy (anonymous):

4+/- (square root of) -8 (over )4

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

\[F (x)= x^3+2x^2-6x=x(x^2+2x-6)\] one zero is 0, the other you get from solving \[x^2+2x-6=0\]

OpenStudy (anonymous):

or are you on the first one

OpenStudy (anonymous):

Either is fine with me, let me try to do what you said....so to solve for x ?

OpenStudy (anonymous):

yes 2 is even so skip the quadratic formula

OpenStudy (anonymous):

easier to complete the square do you don't have to rewrite in simplest radical form and cancel

OpenStudy (anonymous):

it it is totally unclear what i am saying, let me know that is for solving \[x^2+2x-6=0\]

OpenStudy (anonymous):

Would it be +/- (square root of ) 3 ? Or was I not suppose to drop off that other x ?

OpenStudy (anonymous):

i don't think so no

OpenStudy (anonymous):

\[x^2+2x-6=0\\ x^2+2x=6\] is a start

OpenStudy (anonymous):

now what is half of 2?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

and 1 squared?

OpenStudy (anonymous):

Still 1 ?.....

OpenStudy (anonymous):

Unless if negative the i

OpenStudy (anonymous):

yes so go right from \[x^2+2x=6\] so \[(x+1)^2=6+1\\ (x+1)^2=7\]

OpenStudy (anonymous):

that means \[x+1=\pm\sqrt7\\ x=-1\pm\sqrt7\]

OpenStudy (anonymous):

no quadratic formula was hurt in this process you would get the same answer, but it would take a lot more work to get it

OpenStudy (anonymous):

Oh I see, yes.... my professor definitely didn't go over this! And yes this seems simpler

OpenStudy (anonymous):

Thank you ! 😁 @satellite73